3- Solve the following improper integrals 1=e dt 4- Solve the following improper integrals I
Answer #1

3

footvtedt = 3V5 – 3V 7 ert(0) Steps SotVie hdt Hide Steps Compute the indefinite integral: ſt tedt= -e StVtedt Hide Steps t/tShow Steps Apply Integration By Parts: u=ti,v=e-t =-*-5--Vedt Hide Steps s-le-Vid=-(-2Vi + verf ( 17 )) s - e *Vedt Take theApply u-substitution: u=vt Show Steps = S2e=uau Take the constant out: Sa-f(x) dx = a - Sf(x) == 2.5e-up du Use the nonelemeMultiply fractions: abd =2.0. се се = -1,2V erf(17) Cancel the common factor: 2 = -erf(V7 )?. Ya Multiply: 1.VTOVT =-Verf(VT)Simplify =-*+}(-e+Ve + van art(v7)) Add a constant to the solution 6 =-+*+(-eve + sa ( V7 )) +c Compute the boundaries: Soot

4)